# Magnon-induced transparency of a disordered antiferromagnetic Josephson junction

### v1

@@ Abstract
where $\mathcal{G}_{q_{x}R}(\omega)=\sum_{\mathbf{k}\lambda}\mathcal{G}^{\lambda\lambda}_{\mathbf{k}q_{x}R}(\omega)$ and $\sum_{\mathbf{k}\lambda}\mathcal{G}^{\lambda\lambda}_{\mathbf{k}q_{x}R}(\omega)$ is given by Eq.(S5). It is seen from Eqs.(I.1)-(I.1), Eq.(5) and Eq.(S17) that $B_{q_{x}L}^{11}=-B^{22*}_{L}=iC_{q_{x}L}(\omega)\exp(i\varphi)$, where the function $C_{q_{x}L}(\omega)$ is real. Therefore, $j_{L}$ has the typical phase dependence $j=j_{cL}\sin\varphi$, where $j_{cL}=-4e\sum_{q_{x}}(w_{q_{x}})^{2}\int\frac{d\omega}{2\pi}C_{q_{x}L}(\omega)$ is the critical current. The symmetrized critical current $j_{c}$ is given by $j_{c}=(j_{cL}+j_{cR})/2$.
@@
There are several ways to produce coherent magnetic excitations of the Néel order. Here we consider a set up which integrates a spin-Hall antiferromagnetic nanooscillator and the AFM layer in JJ (Fig.S2). The nanooscillator is combined of a heavy metal film which pumps the spin current into a biaxial antiferromagnetic insulator. As a result, the latter enters the autooscillation mode and generates coherent spin oscillations Khymyn ; Cheng . Then, due to spin-spin exchange interaction these oscillations are transmitted to AFM magnetic moments. Most efficiently it occurs when both antiferromagnets make contact through their uncompensated surfaces. The contact interface is assumed to be located in the $x,y$ plane. We consider the oscillator which was proposed in Ref.Khymyn ; Khymyn2 . There, the staggered magnetization is precessing within the $x,y$ easy plane. The precession frequency is controlled by the DC electric current in the heavy metal film. In addition, weaker higher frequency harmonics also take place. The AFM is a uniaxial antiferromagnet whose easy axis is parallel to the $z$-axis. Due to the exchange interaction on the interface the spin dynamics of the nanooscillator leads to precession of AFM localized spins around this axis. Note, that the choice of the antiferromagnetic nanooscillator in the considered set up is important for two reasons: its higher oscillation frequency and the absence of stray magnetic fields near superconducting contacts.
@@
Figure S2: Color online. The suggested set up for studying the effect of magnetic excitations on the Josephson current through AFM.
@@
In order to evaluate the spin precession amplitude $m$ in Eq.(S13), we follow the approach used in Khymyn2 . Let $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ be magnetizations of two sublattices in AFM, such that $M_{1}=M_{2}=M_{s}$. As usual, let us introduce the dimensionless staggered magnetization $\mathbf{l}=(\mathbf{M}_{1}-\mathbf{M}_{2})/2M_{s}$ and magnetization field $\mathbf{m}=(\mathbf{M}_{1}+\mathbf{M}_{2})/2M_{s}$, where $M_{s}$ is the saturation magnetization. Similar fields $\mathbf{l}^{\prime}$ and $\mathbf{m}^{\prime}$ are introduced for sublattice magnetizations in the nanooscillator. The magnetic fields $\mathbf{m}$ and $\mathbf{m}^{\prime}$ are coupled on the uncompensated interface at $z=0$ via the exchange interaction $E_{s}\mathbf{m}\cdot\mathbf{m}^{\prime}$. For small oscillations of $\mathbf{m}$ around the $z$ axis the equation of motion for $\mathbf{l}$ is given by
@@
where $i\in(x,y)$, $c$ is the spin wave velocity and $\Omega_{r}$ is the frequency of the AFM resonance. The solution of this equation has the form $l_{i}=A_{i}\exp(z/\lambda)+B_{i}\exp(-z/\lambda)$, where $\lambda^{-1}=c^{-1}(\Omega_{r}^{2}-\Omega^{2})^{1/2}$. In the considered regime of $\Omega<\Omega_{r}$ $\lambda$ is real and the magnetic excitation in AFM is the evanescent wave. The boundary conditions on the free surface $(z=d)$ and on the interface with the antiferromagnetic insulator $(z=0)$ are given by

### v2

@@ Abstract
where $\mathcal{G}_{q_{x}R}(\omega)=\sum_{\mathbf{k}\lambda}\mathcal{G}^{\lambda\lambda}_{\mathbf{k}q_{x}R}(\omega)$. Its Keldysh component $\mathcal{G}^{K}_{q_{z}R}(\omega)$ is given by Eq.(S5), while retarded and advanced functions $\mathcal{G}^{r(a)}_{q_{z}R}(\omega)$ can be obtained from Eq.(10).
[…]
@@
In the previous section we obtained the expression for the Josephson current which is induced by magnetic excitations. This effect depends on their amplitude $m$ which enters into the electron-magnon coupling constant $F$ given by Eq.(S13). There are several ways to produce coherent magnetic excitations of the Néel order. Here we consider a set up which integrates a spin-Hall antiferromagnetic nanooscillator and the AFM layer in JJ (Fig.S2). The nanooscillator is combined of a heavy metal film which pumps the spin current into a biaxial antiferromagnetic insulator. As a result, the latter enters the autooscillation mode and generates coherent spin oscillations Khymyn ; Cheng . Then, due to spin-spin exchange interaction these oscillations are transmitted to AFM magnetic moments. Most efficiently it occurs when both antiferromagnets make contact through their uncompensated surfaces. The contact interface is assumed to be located in the $x,z$ plane. We consider oscillators which were proposed in Refs.Khymyn ; Cheng . In the high excitation regime the staggered magnetization there precesses within the $x,y$ plane. However, in Ref.Khymyn the latter is the easy plane of the nanooscillator, while in Ref.Cheng it is perpendicular to the easy plane and the easy axis is parallel to the $z$-axis in Fig.S2. The precession frequency is controlled by the DC electric current in the heavy metal film. In addition, weaker higher frequency harmonics also take place. The AFM is a uniaxial antiferromagnet whose easy axis is parallel to the $z$-axis. Due to the exchange interaction on the interface the spin dynamics of the nanooscillator leads to precession of AFM localized spins around this axis. Note, that the choice of the antiferromagnetic nanooscillator in the considered set up is important for two reasons: its higher oscillation frequency and the absence of stray magnetic fields near superconducting contacts.
It is important that ferromagnetic and antiferromagnetic nanooscillators operate at high dc currents in heavy metals. Therefore, Joule heating from such devices must be removed to keep cryogenic conditions in a Josephson junction. At least for the ferromagnetic oscillator such conditions have been provided in Ref.Liu . It is reasonable to expect that such cryogenic conditions can also be realized for antiferromagnetic oscillators.
@@
Figure S2: Color online. The suggested set up for studying the effect of magnetic excitations on the Josephson current through AFM. The electric current in the Pt layer is directed parallel to the $x$ axis. The spin-Hall current is polarized along the $z$ axis
@@
In order to evaluate the spin precession amplitude $m$ in Eq.(S13), we follow the approach used in Khymyn2 . Let $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ be magnetizations of two sublattices in AFM, such that $M_{1}=M_{2}=M_{s}$. As usual, let us introduce the dimensionless staggered magnetization $\mathbf{l}=(\mathbf{M}_{1}-\mathbf{M}_{2})/2M_{s}$ and magnetization field $\mathbf{m}=(\mathbf{M}_{1}+\mathbf{M}_{2})/2M_{s}$, where $M_{s}$ is the saturation magnetization. Similar fields $\mathbf{l}^{\prime}$ and $\mathbf{m}^{\prime}$ are introduced for sublattice magnetizations in the nanooscillator. These fields are coupled on the interface at $y=0$ via the interface exchange interaction $E_{s}(\mathbf{m}+\kappa\mathbf{l})\cdot(\mathbf{m}^{\prime}+\kappa^{\prime}\mathbf{l}^{\prime})$, where $\kappa$ and $\kappa^{\prime}$ are phenomenological parameters. So, when these parameters are equal to 1, both interfaces are uncompensated. At the same time, when they turn to 0 both interfaces are fully compensated. Since for typical antiferromagnets $m\ll l$, the strongest interface exchange interaction takes place for fully uncompensated interfaces.
For small oscillations of $\mathbf{m}$ around the $z$ axis the equation of motion for $\mathbf{l}$ is given by
@@
where $i\in(x,y)$, $c$ is the spin wave velocity and $\Omega_{r}$ is the frequency of the AFM resonance. The Gilbert damping has been neglected in this equation. Such an approximation is valid, as long as the width of the antiferromagnetic resonance, which is caused by this damping, is smaller than the difference between nanooscillator and AFM resonance frequencies. The solution of this equation has the form $l_{i}=A_{i}\exp(z/\lambda)+B_{i}\exp(-z/\lambda)$, where $\lambda^{-1}=c^{-1}(\Omega_{r}^{2}-\Omega^{2})^{1/2}$. In the considered regime of $\Omega<\Omega_{r}$ $\lambda$ is real and the magnetic excitation in AFM is the evanescent wave. The boundary conditions on the free surface $(z=d)$ and on the interface with the antiferromagnetic insulator $(z=0)$ are given by
@@
It is important that the antiferromagnetic insulator NiO was chosen in Refs. Khymyn and Cheng as a part of the nanooscillator. In this case, if the hard axis of NiO is parallel to the $z$ axis in Fig(2S), its interface with AFM is magnetically compensated. Therefore, the device which was considered in Khymyn can only weakly interact with AFM, because $l^{\prime}$ in Eq.(S28) should be substituted for much smaller $m^{\prime}$. On the other hand, in the case considered in Ref.Cheng the uncompensated interface magnetization, whose easy axis is parallel to $z$, might take place.